A Bibliography of Publications of Jingling Xue

Jingling Xue

Department of Mathematics, Statistics and Computing Science

Armidale, NSW 2351 Australia

Tel: +61 67 73 3149 FAX: +61 67 73 3312

E-mail: xue@neumann.une.edu.au (Internet)

18 February 2016 Version 1.06

Abstract

This bibliography records publications of Jingling Xue.

Title word cross-reference

92-VAPP [BCRT92]. 9th [IEE95].

access [Xue92b]. Adapting [LX94]. algorithm [LX94, Xue93a]. Algorithms [LX91, QR91, XL94]. Application [DW93]. Application-Specific [DW93]. Approaches [Meg94]. April [IEE95]. architectures [QR91]. Array [DW93, LX91]. Arrays [XL91, Xue92a, XL92, XL94, Xue91, Xue92b, Xue93b, Xue95a]. automate [Xue93a]. Automating [Xue94]. Avoiding [XL94].

Barbara [IEE95]. Bonas [QR91].

California [IEE95]. Chateau [QR91].

China [Ni94]. Closed [Xue95a]. Closed-form [Xue95a]. Compiling [Xue95b]. Computational [XL94]. conditions [Xue93b, Xue95a]. Conference [BCRT92, DW93, Ni94]. Conflicts [XL94]. CONPAR [BCRT92]. Constructing [Xue95b]. Control [XL91, Xue92a, XL92, Xue91]. Convex [Xue95b].

Dallas [IEE93]. Data [XL94, Xue92b]. December [IEE93, Ni94]. Design [Meg94, LX94]. Dimensional [XL91, XL92, XL94]. Distributed [IEE93, Ni94]. DO [Xue95b].

Equations [XL91, Xue91].

Fifth [IEE93]. form [Xue95a]. Formal [Xue92a]. formulation [Xue93b]. France [BCRT92, QR91].

Gers [QR91].

ICPADS’94 [Ni94]. IEEE [IEE93]. II [QR91]. International [BCRT92, DW93, IEE95, Ni94, QR91]. Italy [DW93]. Iteration [Xue95b].

Joint [BCRT92]. June [QR91].

linear [Xue93b, Xue95a]. Link [XL94]. loading [Xue92b]. Loop [Xue94, Xue93a]. Loops [Xue95b]. Lower [XL94]. Lower-Dimensional [XL94]. Lyon [BCRT92].

Machines [Xue95b]. mapping [Xue93b, XL94, Xue95a]. Massive [Xue94].

nests [Xue93a]. Non [Xue94, Xue95b, Xue93a]. Non-Convex [Xue95b]. Non-unimodular [Xue94, Xue93a].

October [DW93]. One [XL91, XL92]. One-Dimensional [XL91, XL92].

Parallel [BCRT92, IEE93, IEE95, Ni94, Xue95b, QR91]. Parallelism [Xue94]. Proceedings [IEE93, IEE95, Ni94, BCRT92, DW93, QR91]. Processing [IEE93, IEE95, BCRT92]. Processor [XL94, Xue95a]. Processors [DW93]. Pyramidal [LX91].

recovery [Xue92b]. Recurrence [XL91, Xue91]. Republic [Ni94].

Santa [IEE95]. Second [BCRT92]. September [BCRT92]. sequential [LX94]. Signals [XL91, Xue92a, XL92, Xue91]. Spaces [Xue95b]. Specific [DW93]. Specifying [XL91, Xue91]. stationary [Xue92b]. Symposium [IEE93, IEE95]. Synthesis [Xue92a, XL92, Xue93b, Xue95a]. Systems [Ni94]. Systolic [LX91, Meg94, XL91, Xue92a, XL92, LX94, Xue91, Xue92b, Xue93b].

Taiwan [Ni94]. Texas [IEE93]. Transformational [Meg94]. Transformations [Xue94, Xue93a].

Uniform [XL91, Xue91]. unimodular [Xue93a, Xue94].

V [BCRT92]. VAPP [BCRT92]. Vector [BCRT92]. Venice [DW93]. VLSI [QR91].

Workshop [QR91].

References

Bouge:1992:PPC

[BCRT92] L. Bouge, M. Cosnard, Y. Robert, and D. Trystram, editors. Parallel processing: CONPAR 92-VAPP V, Second Joint International Conference on Vector and Parallel Processing, Lyon, France, September 1–4, 1992: proceedings, volume 634 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / New York, NY, USA, etc., 1992. ISBN 3-540-55895-0 (Berlin), 0-387-55895-0 (N.Y.). LCCN QA76.58 .J65 1992.

Dadda:1993:ICA

[DW93] Luigi Dadda and Benjamin W. Wah, editors. The International Conference on Application-Specific Array Processors: October 25– 27, 1993, Venice, Italy: proceedings. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3492-8 (case), 0-8186-3491-X (microfiche). LCCN TK5102.5 .I58 1993. IEEE Computer Society Press order number 3492-02; IEEE catalog number 93TH0572-8.

IEEE:1993:PFI

[IEE93] IEEE, editor. Proceedings of the Fifth IEEE Symposium on Parallel and Distributed Processing: December 1–4, 1993, Dallas, Texas. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4221-1 (microfiche), 0-8186-4222-X (case bound). LCCN QA76.58.I42 1993. IEEE catalog number 93TH0584-3; IEEE Computer Society Press order number 4222-02.

IEEE:1995:PIP

[IEE95] IEEE, editor. Proceedings / 9th International Parallel Processing Symposium, April 25–28, 1995, Santa Barbara, California. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7074-6. LCCN QA76.58 .I58 1995. IEEE Catalog Number: 95TH8052.

Lengauer:1991:SAP

[LX91] C. Lengauer and J. Xue. A systolic array for pyramidal algorithms. Journal of VLSI Signal Processing, 3(3):239–259, 1991. CODEN JVSPED. ISSN 0922-5773 (print), 1573-109x (electronic).

Lengauer:1994:ASA

[LX94] C. Lengauer and J. Xue. Adapting a sequential algorithm for a systolic design. In Megson [Meg94], chapter 8, pages 179–204. ISBN 0-412-44830-0. LCCN QA76.9.A73 T73 1994. US£40.00.

Megson:1994:TAS

[Meg94] G. M. Megson, editor. Transformational Approaches to Systolic Design, volume 2 of Parallel and distributed computing series. Chapman and Hall, Ltd., London, UK, 1994. ISBN 0-412-44830-0. xiv + 301 pp. LCCN QA76.9.A73 T73 1994. US£40.00.

Ni:1994:PII

[Ni94] Lionel M. Ni, editor. Proceedings: ICPADS’94: 1994 International Conference on Parallel and Distributed Systems, December 19–21, 1994, Taiwan, Republic of China. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6555-6. LCCN QA76.58 .I5 1994.

Quinton:1991:APV

[QR91] Patrice Quinton and Yves Robert, editors. Algorithms and parallel VLSI architectures II: proceedings of the International Workshop, Algorithms and Parallel VLSI Architectures II, Chateau de Bonas, Gers, France, June 3–6, 1991. Elsevier North-Holland, Inc., New York, NY, USA, 1991. ISBN 0-444-89153-6. LCCN QA76.58 .I57 1991.

Xue:1991:SCSa

[XL91] J. Xue and C. Lengauer. Specifying control signals for one-dimensional systolic arrays by uniform recurrence equations. In Quinton and Robert [QR91], pages 181– 187. ISBN 0-444-89153-6. LCCN QA76.58 .I57 1991.

Xue:1992:SCSa

[XL92] J. Xue and C. Lengauer. The synthesis of control signals for onedimensional systolic arrays. Integration, the VLSI journal, 14 (1):1–32, November 1992. CO-DEN IVJODL. ISSN 0167-9260 (print), 1872-7522 (elec-

tronic).  URL https://www.
sciencedirect.com/science/article/
pii/016792609290008M.

Xue:1994:ADL

[XL94] J. Xue and P. Lenders. Avoiding data link and computational conflicts in mapping algorithms to lower-dimensional processor arrays. In Ni [Ni94], pages 567–

572. ISBN 0-8186-6555-6. LCCN QA76.58 .I5 1994.

Xue:1991:SCSb

[Xue91] Jingling Xue. Specifying control signals for systolic arrays by uniform recurrence equations. Parallel Processing Letters, 1(2):83–93, December 1991. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Xue:1992:FSC

[Xue92a] J. Xue. The Formal Synthesis of Control Signals for Systolic Arrays. PhD thesis, Department of Computer Science, University of Edinburgh, July 1992.

Xue:1992:LRA

[Xue92b] J. Xue. On the loading, recovery and access of stationary data in systolic arrays. In Bouge et al.

[BCRT92], pages 259–264. ISBN 3-540-55895-0 (Berlin), 0-387-55895-0 (N.Y.). LCCN QA76.58 .J65 1992.

Xue:1993:AAN

[Xue93a] J. Xue. An algorithm to automate non-unimodular transformations of loop nests. In IEEE [IEE93], pages 512–519. ISBN 0-8186-4221-1 (microfiche), 0-8186-4222-X (case bound). LCCN QA76.58.I42 1993. IEEE catalog number 93TH0584-3; IEEE Computer Society Press order number 4222-02.

Xue:1993:NFM

[Xue93b] J. Xue. A new formulation of mapping conditions for the synthesis of linear systolic arrays. In Dadda and Wah [DW93], pages 297– 308. ISBN 0-8186-3492-8 (case), 0-8186-3491-X (microfiche). LCCN TK5102.5 .I58 1993. IEEE Computer Society Press order number 3492-02; IEEE catalog number 93TH0572-8.

Xue:1994:ANU

[Xue94] J. Xue. Automating nonunimodular loop transformations for massive parallelism. Parallel Comput., 20(5):711–728, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Xue:1995:CFM

[Xue95a] J. Xue. Closed-form mapping conditions for the synthesis of linear processor arrays. Journal of VLSI Signal Processing, 10(2):181–199, 1995. CODEN JVSPED. ISSN 0922-5773 (print), 1573-109x (electronic).

Xue:1995:CDL

[Xue95b] J. Xue. Constructing DO loops for non-convex iteration spaces in compiling for parallel machines. In IEEE [IEE95], pages 364– 368. ISBN 0-8186-7074-6. LCCN QA76.58 .I58 1995. IEEE Catalog Number: 95TH8052.